You want to obtain a sample to estimate a population mean. Based on previous evidence, you believe the population standard deviation is approximately σ=53.4σ=53.4. You would like to be 90% confident that your esimate is within 0.2 of the true population mean. How large of a sample size is required?

n =
Answer #1

Solution :

Given that,

Population standard deviation = \sigma = 53.4

Margin of error = E = 0.2

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

sample size = n = (Z\alpha/2* \sigma / E) 2

n = (1.645 * 53.4 / 0.2)2

n = 192909.82

n = 192910

Sample size = 192910